sample of 51 observations will be taken from a process (an infinite population). The population proportion equals 0.85. The probability that the sample proportion will be between 0.9115 and 0.946 is _____.
Answer #1

Solution

Given that,

=  [p ( 1 - p ) / n] = [(0.85 * 0.15) / 51] = 0.05

= P[(0.9115 - 0.85) /0.05 < ( - ) / < (0.946 - 0.85) / 0.05]

= P(1.23 < z < 1.92)

= P(z < 1.92) - P(z < 1.23)

= 0.9726 - 0.8907

= 0.0819

Probability = 0.0819